# Multi-dimensional optimisation of the scanning strategy for the LiteBIRD space mission

### v1

@@ 7.2 Sky pixel visit/revisit times
On the other hand, LiteBIRD and PICO have a larger $\alpha$ and shorter precession period than Planck. This reduces the size of the ring described by the spin and allows the pointing to visit many pixels on the sky many occasions for long periods of time due to its relatively fast precession. A characteristic difference between the two spacecraft is the width of the gap found in the center of the visit time distribution of equatorial pixels. This gap is due to the scan pupil discussed in section 5.1, which discussed now no observations are made while pixels are located within the scan pupil. Note that the size of the scan pupil is determined by $2|\alpha-\beta|$, so there is an annual unobservable time determined by the scan pupil size and the orbital velocity in regions that are $|\alpha-\beta|$ away from the equator.
Figure 12 is a histogram of the revisit time against the visit time in figure 11, showing how many revisit events of what time scale occur over the period of visiting the three considered pixels, with the blue and red dashed lines representing the spin period and precession period, respectively. For the cases of LiteBIRD and PICO, the three pixels can be visited for long periods of time and the wide distribution of revisit times has the potential to construct a rich set of null-tests of various time scales, while for the case of Planck, short timescale revisit event is rich, but the pointing can only visit a specific pixel for a very on period. This restricts null-tests, especially those requiring long timescale data sets.
@@ 7.3 Planet visit/revisit times
A long integration time of planet observations discussed in section 5.1 is one of the important aspects for in-flight calibration, but the frequency and duration of each observation allowing for calibration during the mission is also important. In this section, visit and revisit times are simulated for target planets to clarify the relationship between the scanning strategy and the visibility of the planet. The simulations use the same setup as in section 5.1, except that in that section the visit time was integrated over time, whereas here we consider the visit time distribution before integration.
@@ 8 Discussion and conclusion
In this study, we have explored the parameter space of the scanning strategy of a spacecraft mission equipped with a HWP, using $\{\alpha,T_{\alpha},T_{\beta}\}$ as free parameters. We examined the impact of these parameters on three crucial metrics for $B$-mode observation: visibility time of planets; hit-map uniformity; and cross-link factor. In particularly, the relationship between the cross-link factor and the bias on the tensor-to-scalar ratio, $r$, i.e., $\Delta r$, which has been deemed important for suppressing systematic effects, suggests that the spin-based map-making we introduced in section 4 could serve as an interface for understanding the nontrivial propagation of systematic effects in scanning strategies.
@@ Acknowledgments
This work is supported in Japan by ISAS/JAXA for Pre-Phase A2 studies, by the acceleration program of JAXA research and development directorate, by the World Premier International Research Center Initiative (WPI) of MEXT, by the JSPS Core-to-Core Program of A. Advanced Research Networks, and by JSPS KAKENHI Grant Numbers JP15H05891, JP17H01115, and JP17H01125. The Canadian contribution is supported by the Canadian Space Agency. The French LiteBIRD phase A contribution is supported by the Centre National d’Etudes Spatiale (CNES), by the Centre National de la Recherche Scientifique (CNRS), and by the Commissariat à l’Energie Atomique (CEA). The German participation in LiteBIRD is supported in part by the Excellence Cluster ORIGINS, which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy (Grant No. EXC-2094 - 390783311). The Italian LiteBIRD phase A contribution is supported by the Italian Space Agency (ASI Grants No. 2020-9-HH.0 and 2016-24-H.1-2018), the National Institute for Nuclear Physics (INFN) and the National Institute for Astrophysics (INAF). Norwegian participation in LiteBIRD is supported by the Research Council of Norway (Grant No. 263011) and has received funding from the European Research Council (ERC) under the Horizon 2020 Research and Innovation Programme (Grant agreement No. 772253 and 819478). The Spanish LiteBIRD phase A contribution is supported by MCIN/AEI/10.13039/501100011033, project refs. PID2019-110610RB-C21, PID2020-120514GB-I00, PID2022-139223OB-C21 (funded also by European Union NextGenerationEU/PRTR), and by MCIN/CDTI ICTP20210008 (funded also by EU FEDER funds). Funds that support contributions from Sweden come from the Swedish National Space Agency (SNSA/Rymdstyrelsen) and the Swedish Research Council (Reg. no. 2019-03959). The US contribution is supported by NASA grant no. 80NSSC18K0132. YT acknowledges the support by the Grant-in-Aid for JSPS Fellows, JSPS core-to-core program, JST the establishment of university fellowships towards the creation of science technology innovation, Grant no. JP23KJ1602, JPJSCCA20200003, and JPMJFS2128. GP and HI acknowledge the JSPS Invitation Fellowships for Research in Japan. LV acknowledges partial support by the Italian Space Agency LiteBIRD Project (ASI Grants No. 2020-9-HH.0 and 2016-24-H.1-2018), as well as the RadioForegroundsPlus Project HORIZON-CL4-2023-SPACE-01, GA 101135036.
@@ A.1 Scanning motion of spacecraft
B.2 Discussion on the rotation direction of the spacecraft

### v2

@@ 7.2 Sky pixel visit/revisit times
On the other hand, LiteBIRD and PICO have a larger $\alpha$ and shorter precession period than Planck. This reduces the size of the ring described by the spin and allows the pointing to visit many pixels on the sky many occasions for long periods of time due to its relatively fast precession. A characteristic difference between the two spacecraft is the width of the gap found in the center of the visit time distribution of equatorial pixels. This gap is created by the fact that the equatorial sky pixels are contained within the scan pupil mentioned in section 5.1. Note that the size of the scan pupil is determined by $2|\alpha-\beta|$, so there is an annual unobservable time determined by the scan pupil size and the orbital velocity in regions that are $|\alpha-\beta|$ away from the equator.
Figure 12 is a histogram of the revisit time against the visit time in figure 11, showing how many revisit events of what time scale occur over the period of visiting the three considered pixels, with the blue and red dashed lines representing the spin period and precession period, respectively. For the cases of LiteBIRD and PICO, the three pixels can be visited for long periods of time and the wide distribution of revisit times has the potential to construct a rich set of null-tests of various time scales, while for the case of Planck, short timescale revisit event is rich, but the pointing can only visit a specific pixel for a very short period of time. This restricts null-tests, especially those requiring long timescale data sets.
@@ 7.3 Planet visit/revisit times
The long integration time of planet observations discussed in section 5.1 is one important aspect of in-flight calibration, but the length of time that calibration is possible during the mission and how often calibration can be performed are also important. In this section, visit and revisit times are simulated for target planets to clarify the relationship between the scanning strategy and the visibility of the planet. The simulations use the same setup as in section 5.1, except that in that section the visit time was integrated over time, whereas here we consider the visit time distribution before integration.
@@ 8 Discussion and conclusion
In this study, we have explored the parameter space of the scanning strategy of a spacecraft mission equipped with a HWP, using $\{\alpha,T_{\alpha},T_{\beta}\}$ as free parameters. We examined the impact of these parameters on three crucial metrics for $B$-mode observation: visibility time of planets; hit-map uniformity; and cross-link factor. In particular, the map-based simulations using the spin formalism that we introduced, helps us to understand the relationship between the bias of tensor-to-scalar ratio $r$, or $\Delta r$, and cross-link factors, which have been considered important in order to suppress systematic effects. The coupling between scanning strategies and systematic effects for HWP-equipped spacecraft has been non-trivial, but this method may provide an interface to understand it.
@@ Acknowledgments
This work is supported in Japan by ISAS/JAXA for Pre-Phase A2 studies, by the acceleration program of JAXA research and development directorate, by the World Premier International Research Center Initiative (WPI) of MEXT, by the JSPS Core-to-Core Program of A. Advanced Research Networks, and by JSPS KAKENHI Grant Numbers JP15H05891, JP17H01115, and JP17H01125. The Canadian contribution is supported by the Canadian Space Agency. The French LiteBIRD phase A contribution is supported by the Centre National d’Etudes Spatiale (CNES), by the Centre National de la Recherche Scientifique (CNRS), and by the Commissariat à l’Energie Atomique (CEA). The German participation in LiteBIRD is supported in part by the Excellence Cluster ORIGINS, which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy (Grant No. EXC-2094 - 390783311). The Italian LiteBIRD phase A contribution is supported by the Italian Space Agency (ASI Grants No. 2020-9-HH.0 and 2016-24-H.1-2018), the National Institute for Nuclear Physics (INFN) and the National Institute for Astrophysics (INAF). Norwegian participation in LiteBIRD is supported by the Research Council of Norway (Grant No. 263011) and has received funding from the European Research Council (ERC) under the Horizon 2020 Research and Innovation Programme (Grant agreement No. 772253 and 819478). The Spanish LiteBIRD phase A contribution is supported by MCIN/AEI/10.13039/501100011033, project refs. PID2019-110610RB-C21, PID2020-120514GB-I00, PID2022-139223OB-C21 (funded also by European Union NextGenerationEU/PRTR), and by MCIN/CDTI ICTP20210008 (funded also by EU FEDER funds). Funds that support contributions from Sweden come from the Swedish National Space Agency (SNSA/Rymdstyrelsen) and the Swedish Research Council (Reg. no. 2019-03959). The US contribution is supported by NASA grant no. 80NSSC18K0132. This work has also received funding by the European Union’s Horizon 2020 research and innovation program under grant agreement no. 101007633 CMB-Inflate. YT acknowledges the support by the Grant-in-Aid for JSPS Fellows, JSPS core-to-core program, JST the establishment of university fellowships towards the creation of science technology innovation, Grant no. JP23KJ1602, JPJSCCA20200003, and JPMJFS2128. GP and HI acknowledge the JSPS Invitation Fellowships for Research in Japan. LV acknowledges partial support by the Italian Space Agency LiteBIRD Project (ASI Grants No. 2020-9-HH.0 and 2016-24-H.1-2018), as well as the RadioForegroundsPlus Project HORIZON-CL4-2023-SPACE-01, GA 101135036.
@@ A.1 Scanning motion of spacecraft
B.2 Metrics for detectors located away from the boresight
In this section we investigate how the metrics of detectors located away from the boresight (i.e. center of focal plane) behave in our standard configuration. Different detectors on the focal plane are often described as having different effective $\beta$ (written as $\beta^{\rm eff}$) than the one of the boresight. This can be expressed by considering the shifted effective spin axes per detector as
$\beta^{\rm eff}_{i}=\beta+\beta_{i},$ (B.1)
where $i$ represents detector index. Detectors that observe in a direction that is inclined by $\beta_{i}$ from the boresight have effectively different geometric parameters, so the scan trajectory is different from that of the boresight and, in general, the metrics also have different values. Conversely, a detector that is not boresight but whose $\beta^{\rm eff}_{i}$ is close to $\beta$ has a trajectory that is almost redundant with the boresight, so the metrics do not change much. In the MFT/HFT field of view with $14^{\circ}$ in radius, we simulated the integration time of planet visibility and cross-link factor by a detector which has minimum/maximum $\beta_{i}$ ($-14^{\circ}/14^{\circ}$).
Section B.2 shows integration time of planets visibility (top left: $\beta_{i}=-14^{\circ}$, top middle: $\beta_{i}=14^{\circ}$). As in figure 3 (left), the results for the integration time of planets visibility are independent of $T_{\alpha}$, with a peak at $\alpha=(\kappa+\beta_{i})/2$. From section B.2 (top left/middle), it is clear that our standard configuration is a scanning strategy that provides relatively long integration time of planet visibility not only at the boresight but also at the detector with the minimum or maximum $\beta^{\rm eff}_{i}$.
Section B.2 shows spin-$(2,0)$ cross-link factor (top right: $\beta_{i}=-14^{\circ}$, bottom left: $\beta_{i}=14^{\circ}$), and spin-$(2,4)$ cross-link factor (bottom middle: $\beta_{i}=-14^{\circ}$, bottom right: $\beta_{i}=14^{\circ}$). When considering the effective $\beta$, certain detectors cannot satisfy eq. 2.4, and the area around the pole of the ecliptic coordinates becomes unobservable, so the cross-links factor in that region diverge. Here we show the results of calculating the full-sky average using the detector which has $\beta_{i}=\pm 14$, ignoring the pixels where the cross-link factor diverge. First, the $\beta_{i}=-14^{\circ}$ detector is not able to observe the full-sky because its trajectory does not go beyond the poles of the ecliptic coordinates. Comparing the cross-link factors for a detector which has $\beta_{i}=\pm 14^{\circ}$, we can see that the distribution of spin-$(2,0)$ cross-link factor is similar in structure, but the overall value is larger for $\beta_{i}=-14^{\circ}$. This is due to the fact that the uniformity of the crossing angle for each pixel has decreased due to the narrower area which can be accessed by the scanning. The spin-$(2,4)$ cross-link factor also shows a change in structure, and the values are generally smaller for $\beta_{i}=-14^{\circ}$. This is due to the fact that a smaller $\beta_{i}$ slows down the scan speed defined by LABEL:eq:sweeping_velocity, causing the HWP to rotate extra within the sky pixel. It is clear that although the structure and value of the distribution changed, the standard configuration still provides small cross-link factors for detectors other than the boresight.
[Uncaptioned image][Uncaptioned image][Uncaptioned image]
Figure 17: (top left/top middle) Integrated planet visibility time simulated by a detector which has $\beta_{i}=-14^{\circ}/14^{\circ}$. (top right/bottom left) spin-$(2,0)$ cross-link factor simulated by a detector which has $\beta_{i}=-14^{\circ}/14^{\circ}$. (bottom middle/bottom right) spin-$(2,4)$ cross-link factor simulated by a detector which has $\beta_{i}=-14^{\circ}/14^{\circ}$. Values for sky pixels with diverging cross-link factors are ignored and averaged over the entire sky.
B.3 Discussion on the rotation direction of the spacecraft